OpenStudy (anonymous):

Add x+1/2x + 5x/2x+1 Subtract x/x+3 -5x/2x+1 Multiply x-4/x+1 * 5x/2x+1 Divide 6/x+3 / 5x/2x+1 please help

OpenStudy (lasttccasey):

Is this correct for the first one? Add:\[x+\frac{1} {2}x + \frac{5x}{2x}+1\]

OpenStudy (anonymous):

\[\frac{ x+1 }{2x } +\frac{ 5 }{ 2x+1 }\]

OpenStudy (lasttccasey):

Well I don't believe they can be simply "added". Maybe factored?

OpenStudy (anonymous):

yeah i think it needs to be factored and whatnot but i dont know how to do that

OpenStudy (lasttccasey):

Well the (x+1)/(2x) can be simplified by simply splitting this into two fractions. x/2x + 1/(2x) + ... 1/2 + 1/(2x) + ... and the other part is kinda stuck because you can't pull apart the bottom without an equals sign but you can divide top and bottom by two so I'd say 1/2 + 1/(2x) + 2.5/(x+.5)

OpenStudy (anonymous):

is that it?

OpenStudy (lasttccasey):

For the first one. Maybe @SolomonZelman will have more insight for you.